OpenStudy (anonymous):

Find the exact value of tan 13π/8 using half-angle identities. sqrt(2)-sqrt(3)/2 -sqrt(2)-1 1-sqrt(2) sqrt(2)-sqrt(2)/2 possible answers

OpenStudy (anonymous):

i guess \(\frac{13\pi}{8}\) is half of \(\frac{13\pi}{4}\)

OpenStudy (anonymous):

sqrt(2)-sqrt(3)/2 -sqrt(2)-1 1-sqrt(2) sqrt(2)-sqrt(2)/2 possible answers

OpenStudy (anonymous):

oh ok so that half 13x/4?

OpenStudy (anonymous):

perhaps easiest to use \[\tan(\frac{13\pi}{8})=\frac{\sin(\frac{13\pi}{4})}{1+\cos(\frac{13\pi}{4})}\]

OpenStudy (anonymous):

yes \(\frac{13\pi}{8}\) is half of \(\frac{13\pi}{4}\)

OpenStudy (anonymous):

so what the next step to getting the possible answer i listed

OpenStudy (anonymous):

sqrt(2)-sqrt(3)/2 -sqrt(2)-1 1-sqrt(2) sqrt(2)-sqrt(2)/2

OpenStudy (anonymous):

the last step is to compute \(\frac{\sin(\frac{13\pi}{4})}{1+\cos(\frac{13\pi}{4})}\) and see what you get

OpenStudy (anonymous):

i got 1.16 rounded but it doesnt match the other answes so is there something im missing

OpenStudy (anonymous):

since \(\frac{13\pi}{4}\) is coterminal with \(\frac{3\pi}{4}\) you can find \(\sin(\frac{3\pi}{4})\) and \(\cos(\frac{3\pi}{4})\) instead don't use a calculator, write down the exact answer

OpenStudy (anonymous):

oh ok soo use sin=135 and cos 135 too so -sqrt(2)/2 sqrt(2)

OpenStudy (anonymous):

oh ok soo use sin=135 and cos 135 too so -sqrt(2)/2 sqrt(2)/2

OpenStudy (raden):

13pi/4 - 2pi = 5pi/4 i think u can use 5pi/4 or 225 degrees not 135 d...

OpenStudy (anonymous):

oh so use 225 so what do i do next?

OpenStudy (anonymous):

oh damn you are right sorry

OpenStudy (anonymous):

it is coterminal with \(\frac{5\pi}{4}\)

OpenStudy (anonymous):

forget about degrees, you are an adult and so you should only use them if you are "solving" a triangle you need to know that \(\cos(\frac{5\pi}{4})=-\frac{\sqrt{2}}{2}\) and also \(\sin(\frac{5\pi}{4})=-\frac{\sqrt{2}}{2}\)

OpenStudy (anonymous):

so they answer but it doesnt match any of my answers ?? so idk

OpenStudy (anonymous):

\[-\frac{\frac{\sqrt{2}}{2}}{1-\frac{\sqrt{2}}{2}}\]

OpenStudy (anonymous):

simplify by multiplying top and bottom by 2

OpenStudy (anonymous):

\[\frac{-\sqrt{2}}{2-\sqrt{2}}\]

OpenStudy (anonymous):

the answer 2? if u simplfy

OpenStudy (anonymous):

multiply by conyugate of denominator...

OpenStudy (anonymous):

huh so multpy by sqrt 2

OpenStudy (anonymous):

no, but multiply (2+sqrt(2))/(2+sqrt(2)

OpenStudy (anonymous):

1 right

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

???

OpenStudy (anonymous):

can u simplify it :